OpenStudy (anonymous):

The equation (x + 5) + (y + 3) = 169 models the position and range of the source of a radio signal. Describe the position of the source and the range of the signals.

OpenStudy (anonymous):

i am going to make a guess that this is \[(x+5)^2+(y+3)^2=169\] a circle with center \((-5,-3)\) and radius \(\sqrt{169}=13\)

OpenStudy (anonymous):

other wise the question does not really make sense

OpenStudy (anonymous):

that's how the question goes on my paper

OpenStudy (anonymous):

with the square or without the square?

OpenStudy (anonymous):

without

OpenStudy (anonymous):

then i think it was a typo

OpenStudy (anonymous):

it's not a typo

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i will post it again

OpenStudy (anonymous):

The equation (x + 5) + (y + 3) = 169 models the position and range of the source of a radio signal. Describe the position of the source and the range of the signals.

OpenStudy (anonymous):

i think that in this question if i square both sides and then i try to solve the general equation of second degree so formed we can get the solution @satellite73

OpenStudy (anonymous):

?